An amount Q of heat is added to a monatomic ideal gas in a process in which the gas performs a work Q/2 on its surrounding. Find the molar heat capacity for the process.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. Q = Amt of heat given Work done = Q/2, ∆Q = W + ∆ U For monoatomic gas ⇒ ∆U = Q – Q/2 = Q/2 V = n3/2RT = Q/2 = nT * 3/2R = 3R * nT 3RnT = ndTC base P ⇒ C base P = 3R